INFRARED SPECTROSCOPY OF ISOTOPICALLY SUBSTITUTED NO_{3} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

A sufficient concentration of NO_{3} trapped in solid neon has been produced for detection of all of its previously reported infrared bands between 700 and 2600 cm^{−1}, most of which lie within 2.5 cm^{−1} of the gas-phase band origins. Analysis of the infrared spectra obtained for isotopologues of NO_{3} which contain nitrogen-15 and/or oxygen-18 is consistent with a planar, threefold symmetric structure for the ground-state molecule and provides information supporting the occurrence of extensive mixing of ground-state vibrational levels of E′ symmetry and vibronic interaction with the B̃ ^{2}E′ excited electronic state.